# Arrangements of circles supported by small chords and compatible with natural real algebraic functions

### v1

@@ Abstract.
      Let $x_{j^{\prime},0}$ be a point such that the sign of the 1st component and that of the 3rd component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are same and that the sign of the 2nd component and that of the 4th component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are same. Then we can have all cases of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Theorem 1 (2b) and (2c).
@@
      Let $x_{j^{\prime},0}$ be a point such that the sign of the 1st component and that of the 3rd component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are same and that the sign of the 2nd component and that of the 4th component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are mutually distinct. Then we can have all cases of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Theorem 1 (2c) and we cannot have chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Theorem 1 (2b) with $p_{1,2,2}=0$.
@@
      Let $x_{j^{\prime},0}$ be a point such that the sign of the 1st component and that of the 3rd component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are distinct and that the sign of the 2nd component and that of the 4th component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are same. Then we can have all cases of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Theorem 1 (2b) and we cannot have chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Theorem 1 (2c) with $p_{1,2,1}=0$.
@@
      Let $x_{j^{\prime},0}$ be a point such that the sign of the 1st component and that of the 3rd component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are distinct and that the sign of the 2nd component and that of the 4th component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are distinct. Then we can have a case of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Theorem 1 (2a) and this is the unique case.
@@
    Let $x_{j^{\prime},0}$ be a point such that for some straight line $L_{x_{j^{\prime},0}}$, the intersection satisfies $\overline{D_{\mathcal{S}}}\bigcap L_{x_{j^{\prime},0}}=\{x_{j^{\prime},0}\}$. Then we can find chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ to do a discussion similar to that in Theorem 1 (3) with (3a). Here we can also define similar corresponding notions and abuse the terminologies, notions and notation, for example.
@@
We give notes on our proof of Theorem 3. As commented, we can prove Theorem 3 by investigating carefully one by one. We only present FIGUREs 5 and 6 for some cases of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$, for example. ∎
@@
The case (1) is shown by investigating carefully as ever. We only note that in the case the chord at $(D_{\mathcal{S}},x_{j^{\prime},0})$ is as in Theorem 1 (2b) with $p_{1,2,2}=0$, we can argue as Theorem (2a). We have not encountered cases similar to this case and this must be considered in the case Theorem 3 (3a) with (4). See also FIGURE 5.
@@
The local change of the Poincaré-Reeb V-digraph of $D_{\mathcal{S}}$ for ${\pi}_{2,1,1}$ to that of $D_{{\mathcal{S}}^{\prime}}$ for ${\pi}_{2,1,1}$ in the situation of Theorem 1 (2c) and (3b) is same as the case of [7, Theorem 5 with Problem 1] where for the integers ”$i(v_{1})=i(v_{2})$ there” $i(v_{1})>i(v_{2})$ or $i(v_{1})<i(v_{2})$ may hold instead: we can have all cases for the relation between $i(v_{1})$ and $i(v_{2})$.
@@
For the graphs we need, we can understand the definition by checking [7]. More precisely, we only need to check around [7, Theorem 5 with Problem 1] where we do not assume non-trivial mathematical arguments in the definition and formulation. FIGURE 4 shows an example for the case $i(v_{1})>i(v_{2})$: we can have the cases $i(v_{1})<i(v_{2})$ and $i(v_{1})=i(v_{2})$ by considering chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$, according to Theorem 1 (2c) with (3b). ∎
@@
Another simplest example of SSC-NI $(\mathcal{S},D_{\mathcal{S}})$ is shown in FIGURE 7. The set $\mathcal{S}$ consists of exactly $3$ circles where the initial set ${\mathcal{S}}_{0}$ of circles, defined in Definition 3, consists of just one circle: the circle in the left is added according to Theorem 1 (2a) and after that another circle, depicted partially by an arc, is added according to Theorem 3 (3d) with (4).
@@
We close our paper by presenting related problems. Similar problems are also in [7].
@@
Find and formulate other important explicit classes of NI arrangements of circles. Investigate the classes from the viewpoint of combinatorics, singularity theory of smooth functions and maps and real algebraic geometry, for example.

### v2

@@ Abstract.
      Let $x_{j^{\prime},0}$ be a point such that the sign of the 1st component and that of the 3rd component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are same and that the sign of the 2nd component and that of the 4th component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are same. Then we can have all cases of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Proposition 2 (2) and (3).
@@
      Let $x_{j^{\prime},0}$ be a point such that the sign of the 1st component and that of the 3rd component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are same and that the sign of the 2nd component and that of the 4th component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are mutually distinct. Then we can have all cases of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Proposition 2 (3) and we cannot have chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Proposition 2 (2) with $p_{1,2,2}=0$.
@@
      Let $x_{j^{\prime},0}$ be a point such that the sign of the 1st component and that of the 3rd component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are distinct and that the sign of the 2nd component and that of the 4th component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are same. Then we can have all cases of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Proposition 2 (2) and we cannot have chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Proposition 2 (3) with $p_{1,2,1}=0$.
@@
      Let $x_{j^{\prime},0}$ be a point such that the sign of the 1st component and that of the 3rd component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are distinct and that the sign of the 2nd component and that of the 4th component of the signs at $(D_{\mathcal{S}},x_{j^{\prime},0})$ are distinct. Then we can have a case of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ as in Proposition 2 (1) and this is the unique case.
@@
    Let $x_{j^{\prime},0}$ be a point such that for some straight line $L_{x_{j^{\prime},0}}$, the intersection satisfies $\overline{D_{\mathcal{S}}}\bigcap L_{x_{j^{\prime},0}}=\{x_{j^{\prime},0}\}$. Then we can find chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$ to do a discussion similar to that in Theorem 1 (1).
@@
We give notes on our proof of Theorem 3. As the previous theorems, we can prove Theorem 3 by investigating carefully one by one. We only present FIGUREs 5 and 6 for some cases of chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$, for example.
@@
The case (1) is shown by investigating carefully as ever. We only note that in the case the chord at $(D_{\mathcal{S}},x_{j^{\prime},0})$ is as in Proposition 2 (2) with $p_{1,2,2}=0$, we can argue as Proposition 2 (1). We have not encountered cases similar to this case and this must be considered in the case Theorem 3 (1a) with (2). See also FIGURE 5.
@@
The local change of the Poincaré-Reeb V-digraph of $D_{\mathcal{S}}$ for ${\pi}_{2,1,1}$ to that of $D_{{\mathcal{S}}^{\prime}}$ for ${\pi}_{2,1,1}$ in the situation of Proposition 2 (3) with Theorem 1 (2) is same as the case of [7, Theorem 5 with Problem 1] where for the numbers ”$i(v_{1})=i(v_{2})$ there”, sufficiently close to ${\pi}_{2,1,1}(x_{j^{\prime},0})$, the relation $i(v_{1})>i(v_{2})$ or $i(v_{1})<i(v_{2})$ may hold instead. We can have all cases for the relation between the numbers $i(v_{1})$ and $i(v_{2})$, which are sufficiently close to ${\pi}_{2,1,1}(x_{j^{\prime},0})$ and both of which are either smaller or larger than ${\pi}_{2,1,1}(x_{j^{\prime},0})$.
@@
For the graphs we need, we can understand the definition by checking [7]. More precisely, we only need to check around [7, Theorem 5 with Problem 1] where we do not assume non-trivial mathematical arguments in the definition and formulation.
However, we give our related exposition essentially same as the original one.
Let $x_{j^{\prime},0}$ be a vertical pole where the restriction of ${\pi}_{2,1}$ to the boundary of $\overline{D_{\mathcal{S}}}-D_{\mathcal{S}}$ has a local maximum.
There exist exactly two edges $e_{x_{j^{\prime},0},1,-}$ and $e_{x_{j^{\prime},0},2,-}$ which enter $q_{D_{\mathcal{S}},1}(x_{j^{\prime},0})$ and the preimages of the closures of which for the quotient map contain $x_{j^{\prime},0}$. There exist exactly one edge $e_{x_{j^{\prime},0},+}$ which departs from $q_{D_{\mathcal{S}},1}(x_{j^{\prime},0})$ and the preimage of the closure of which for the quotient map contains $x_{j^{\prime},0}$. For the remaining edges entering or departing from $q_{D_{\mathcal{S}},1}(x_{j^{\prime},0})$, we change the vertex they enter or depart from. For each of the two edges $e_{x_{j^{\prime},0},i,-}$ ($i=1,2$), we add two adjacent vertices $v_{x_{j^{\prime},0},i,-,0}$ and $v_{x_{j^{\prime},0},i,-}$ in the interior and sufficiently close to $q_{D_{\mathcal{S}},1}(x_{j^{\prime},0})$. We also define the vertices in such a way that $v_{x_{j^{\prime},0},i,-,0}$ is closer to $q_{D_{\mathcal{S}},1}(x_{j^{\prime},0})$ than $v_{x_{j^{\prime},0},i,-}$. For the remaining edges entering $q_{D_{\mathcal{S}},1}(x_{j^{\prime},0})$, we change the vertex they enter to $v_{x_{j^{\prime},0},1,-,0}$ or $v_{x_{j^{\prime},0},2,-,0}$, according to the location of the preimages for the quotient map, in ${\mathbb{R}}^{2}$. For the edges departing from $q_{D_{\mathcal{S}},1}(x_{j^{\prime},0})$ except $e_{x_{j^{\prime},0},+}$, we change the vertex they depart from to $v_{x_{j^{\prime},0},1,-,0}$ or $v_{x_{j^{\prime},0},2,-,0}$, according to the location of the preimages for the quotient map, in ${\mathbb{R}}^{2}$.
In addition, respecting the structures of our digraphs, we define the structure of the new V-digraph $G_{D_{{\mathcal{S}}^{\prime},1}}$ as the digraph associated with a new function $l_{G_{D_{{\mathcal{S}}^{\prime},1}}}:=V_{D_{{\mathcal{S}}^{\prime}},1}$ on the vertex set whose values at $v_{x_{j^{\prime},0},i,-,0}$ are ${\pi}_{2,1,1}(x_{j^{\prime},0})$, whose values at $v_{x_{j^{\prime},0},i,-}$ are $i(v_{i})<{\pi}_{2,1,1}(x_{j^{\prime},0})$, whose value at $q_{D_{\mathcal{S}},1}(x_{j^{\prime},0})$ is changed into a real number greater than ${\pi}_{2,1,1}(x_{j^{\prime},0})$ and sufficiently close to it, and whose values at the remaining vertices are same as those of the (originally associated) function $l_{G_{D_{\mathcal{S},1}}}:=V_{D_{{\mathcal{S}}},1}$.
FIGURE 4 shows an example for the case $i(v_{1})>i(v_{2})$: we can have the cases $i(v_{1})<i(v_{2})$ and $i(v_{1})=i(v_{2})$ by considering chords at $(D_{\mathcal{S}},x_{j^{\prime},0})$, according to Proposition 2 (3) with Theorem 1 (2).
Let $x_{j^{\prime},0}$ be a vertical pole where the restriction of ${\pi}_{2,1}$ to the boundary of $\overline{D_{\mathcal{S}}}-D_{\mathcal{S}}$ has a local minimum. In this case, we can argue in the same way.
This completes our exposition on the proof. ∎
@@
Another simplest example $(\mathcal{S},D_{\mathcal{S}})$ of SSC-NI is shown in FIGURE 7. The set $\mathcal{S}$ consists of exactly $3$ circles where the initial set ${\mathcal{S}}_{0}$ of circles, defined in Definition 3, consists of just one circle: the circle in the left is added according to Proposition 2 (1) and after that another circle, depicted partially by an arc, is added according to Theorem 3 (1d) with (2).
@@
We close our paper by presenting related problems. Similar problems are also presented in [7].
@@
Find and formulate other important explicit classes of NI arrangements of circles. Investigate the classes from the viewpoint of combinatorics, singularity theory of smooth functions and maps, and real algebraic geometry, for example.